How do I get the answer 1797 using numbers 86254 only once each? You can subtract, add, divide or times in any order.
Answer by greenestamps(13200)



Many (most?) problems like this are a waste of time, because they teach no problem solving skills. But with this one you can get some good mental exercise by doing some logical reasoning to help find a solution. It still requires some trial and error; but the logical reasoning limits the possibilities enough that the amount of trial and error is small.

All the digits are even except the 5; and the last digit of the number you are trying to make is 7. That means the last thing we need to do to get 1797 is either add or subtract 5.

1797+5 = 1802; 1797-5 = 1792. So the problem now is to make either 1802 or 1792 using the digits 2468.

The solution is almost certain to involve multiplication to get a number that large. The product of the 4 digits is too small, so you probably need to make a couple of 2-digit numbers using the 4 remaining digits, and you want the product of those two 2-digit numbers to be either 1802 or 1792.

1792 has far more divisors than 1802; and some trial and error finds 28*64 = 1792.

So...

ANSWER: 28*64+5 = 1797